OpenStudy (anonymous):

find the difference 6rs-(*2rs) explain.

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@Data_LG2

geerky42 (geerky42):

6rs-(*2rs)? what does * stand for?

OpenStudy (anonymous):

they are similar terms, so you can just easily apply the operation but if the "*" has a meaning then maybe the process would be different

OpenStudy (anonymous):

3x^2y * x^2y A.2 B.2x^2y C.3x^2y D. 4x^2y

OpenStudy (anonymous):

i need to find the difference but sometimes a multiplying sign means add

OpenStudy (anonymous):

like i think it would be c but how do i know when i seperate the letter or not?

OpenStudy (mathmath333):

\[3x^{2}y \times x^2y ...?\]

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

like do i add 3 to x? idk

OpenStudy (mathmath333):

Either the options are wrong or the question.

OpenStudy (anonymous):

no this is what it says

OpenStudy (mathmath333):

\[3x^{2}y \times x^2y\\=3x^{2+2}y^{1+1}\\=3x^4y^2\]

OpenStudy (anonymous):

theres a 3x^2y lol

OpenStudy (mathmath333):

typo

geerky42 (geerky42):

"sometimes a multiplying sign means add" So he probably means \[3x^{2}y + x^2y ...?\]

OpenStudy (mathmath333):

i dont think multiplying sign means add sometimes. it must be an error \[\huge \times\implies *\implies .\]

geerky42 (geerky42):

I didn't said I agree with what he said. What I saying is that he may mean \(3x^{2}y + x^2y\) There is correct option for that.